My overall grade is a 80B. Our final weighs 30% . Whats the lowest I can make and still pass with a C 79-70?

Answer by Alan3354(69443)

My overall grade is a 80B. Our final weighs 30% . Whats the lowest I can make and still pass with a C 79-70?
=====================
The overall grade is 70%
70*80 + 30*Final = 100*70 to make 70.
Final = (7000 - 5600)/30 = 46.67%
============
If you make 100 on the final:
70*80 + 30*100 = 100*Grade
Grade = 86%